OpenStudy (anonymous):

A snail moved forward 4/7 inches during one hour, and 1/8 inches during the next hour. How much distance did it cover in the two hours?

OpenStudy (lgbasallote):

wouldn't it just be \[\frac{4}{7} + \frac{1}{8}?\]

OpenStudy (anonymous):

whats the answer though"?

OpenStudy (lgbasallote):

do you know how to add \[\frac 47 + \frac 18?\]

OpenStudy (anonymous):

no

OpenStudy (lgbasallote):

start by making them common denominators..do you know how?

OpenStudy (anonymous):

no